OpenStudy (anonymous):

Explain the difference between a savings account earning simple interest and a savings account earning compound interest.

OpenStudy (anonymous):

@dumbcow @kropot72

OpenStudy (dmezzullo):

With earing compound interest i believe u earn more money than u would with just simple interest

OpenStudy (dmezzullo):

Is that right @dumbcow

OpenStudy (dumbcow):

yes that is true...with simple interest you earn a constant amount of interest each year with compound int you earn a constant rate....the amount increases each year as the balance increases

OpenStudy (anonymous):

so then it would be better to open up a compund interest account right?

OpenStudy (dmezzullo):

If u want more money then yes

OpenStudy (dumbcow):

say you have $100 at 5% interest with simple int you earn $5 per year with compound you earn $5 1st year, then (.05*105) = $5.25 the 2nd year and (.05*110.25) = $5.51 the 3rd year ...

OpenStudy (kobi):

say you have $100 at 5% interest with simple int you earn $5 per year with compound you earn $5 1st year, then (.05*105) = $5.25 the 2nd year and (.05*110.25) = $5.51 the 3rd year ...